1. The equilibrium constant, Kc, for the following reaction is 42.2 at 288 K.

2CH2Cl2(g) <-----> CH4(g) + CCl4(g)  

When a sufficiently large sample of CH2Cl2(g) is introduced into an evacuated vessel at 288 K, the equilibrium concentration of CCl4(g) is found to be 0.140 M.  

Calculate the concentration of CH2Cl2 in the equilibrium mixture. _____ M

2. A student ran the following reaction in the laboratory at 295 K:

2CH2Cl2(g) <-------> CH4(g) + CCl4(g)  

When she introduced 8.46×10-2 moles of CH2Cl2(g) into a 1.00 liter container, she found the equilibrium concentration of CCl4(g) to be 3.90×10-2M.  

Calculate the equilibrium constant, Kc, she obtained for this reaction.  

Kc =
Answer #1

1. At equilibrium , [CH4 ] = [CCl4 ] = 0.140 M

and Kc = 42.2

For given reaction

Kc = [CH4][CCl4] / [CH2Cl2]2

42.2 . = (0.140)2 / [CH2Cl2]2

[CH2Cl2] = 0.0216 M

2. At equilibrium , [CH4 ] = [CCl4 ] = 3.90 *10-2 M = 0.039 M

Initial Moles of CH2Cl2 = 8.46 *10-2 moles = 0.0846 moles

Volume = 1L

so, initial [CH2Cl2 ] = moles / volume = 0.0846 moles/ 1 L = 0.0846 M

[CH2Cl2] reacted = 2 * [CCl4] = 2* 0.039 = 0.078 M

At equilibrium , [CH2Cl2] = Initial value - decomposed value = 0.0846 - 0.078 = 0.0066 M

Now, Kc = [CH4][CCl4] / [CH2Cl2]2

Kc = (0.039)2 / (0.0066)2 = 34.92